# On quasi-Albanese morphisms for log canonical Calabi-Yau pairs

### v1

@@ 2.2. Quasi-Albanese morphism
8Construction.
We recall the constructions of quasi-Albanese morphisms in [Iit76, Fuj24]. Let $(X,D_{X}=\sum_{i=1}^{m}D_{i})$ be a projective log smooth pair with $D_{X}$ reduced. Where $D_{i}$’s are prime divisors. Let $X^{0}:=X\setminus D_{X}$ and let $\iota:X^{0}\to X$ be the open embedding. Let $p:X\to A$ be the Albanese morphism. Set $q=q(X)$, $\overline{q}=q(X,D_{X})$, and $d=\overline{q}-q$.
Let $\eta_{i}$, $1\leq i\leq d$, be a basis of the free part of $\ker(\iota_{*}:H_{1}(X^{0},\mathbb{Z})\to H_{1}(X,\mathbb{Z}))$, and let $\xi_{i}\in H_{1}(X^{0},\mathbb{Z})$, $1\leq i\leq 2q$, be such that these $\eta_{i}$’s and $\xi_{i}$’s form a basis of the free part of $H_{1}(X^{0},\mathbb{Z})$.
Let $\omega_{i}$, $1\leq i\leq q$, be a basis of $H^{0}(\Omega_{X}^{1})$, and let $\phi_{i}\in H^{0}(\Omega_{X}^{1}(\log D))$, $1\leq i\leq d$, be such that these $\omega_{i}$’s and $\phi_{i}$’s form a basis of $H^{0}(\Omega_{X}^{1}(\log D_{X}))$, and satisfy $\int_{\eta_{i}}\omega_{j}=0$ and $\int_{\eta_{i}}\phi_{j}=\delta_{ij}$ (the Kronecker delta).
@@
$A_{i}:=(\int_{\xi_{i}}\omega_{1},\cdots,\int_{\xi_{i}}\omega_{q},\int_{\xi_{i}}\phi_{1},\cdots,\int_{\xi_{i}}\phi_{d})=:(\hat{A}_{i},C_{i}),1\leq i\leq 2q,\text{ and }$
$B_{i}:=(\int_{\eta_{i}}\omega_{1},\cdots,\int_{\eta_{i}}\omega_{q},\int_{\eta_{i}}\phi_{1},\cdots,\int_{\eta_{i}}\phi_{d})=:(0,\hat{B}_{i}),1\leq i\leq d.$
@@
$G:=\mathbb{C}^{\overline{q}}/\sum_{i}\mathbb{Z}A_{i}+\sum_{j}\mathbb{Z}B_{j}.$
@@
$0\to\mathbb{C}^{d}/\sum_{j}\mathbb{Z}\hat{B}_{j}\cong(\mathbb{C}^{*})^{d}\to G\to\mathbb{C}^{q}/\sum_{i}\mathbb{Z}\hat{A}_{i}=A\to 0,$
and $A$ coincides with the Albanese variety of $X$. The isomorphism $\mathbb{C}^{d}/\sum_{j}\mathbb{Z}\hat{B}_{j}\cong(\mathbb{C}^{*})^{d}$ is given by
@@
Remark 9.
Keep the notations introduced in the previous construction. Let $G^{\prime}\to\mathbb{C}^{q}$ be the pullback of the $(\mathbb{C}^{*})^{d}$-bundle $G\to A$ by the cover $\mathbb{C}^{q}\to A$. Then we have $G^{\prime}=\mathbb{C}^{\overline{q}}/\sum_{i}\mathbb{Z}B_{i}\cong\mathbb{C}^{q}\times(\mathbb{C}^{*})^{d}$ and
@@
For each $j$, let $L_{j}\in\mathrm{Pic}^{0}(A)$ be the line bundle corresponding to the representation
(1) $\pi_{1}(A)\to\mathbb{C}^{*},\hat{A}_{i}\mapsto\exp(2\pi\mathrm{i}\int_{\xi_{i}}\phi_{j}),1\leq i\leq q.$
Then $G$ is the fiber product over $A$ of the total spaces of $L_{j}$’s with zero sections removed, in other words,
$G=\operatorname{Spec}_{A}\bigoplus_{(m_{1},\cdots,m_{d})\in\mathbb{Z}^{d}}L_{1}^{-m_{1}}\otimes\cdots\otimes L_{d}^{-m_{d}}.$
$G$ has the following natural compactification:
(2) $G\subset\mathbb{P}_{A}:=\mathbb{P}_{A}(O_{A}\oplus L_{1}^{-1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}^{-1}).$
@@
(2) By the expression $\exp(2\pi\mathrm{i}\int_{\gamma}\phi_{j})$, we mean $\exp(2\pi\mathrm{i}\int_{\gamma^{\prime}}\phi_{j})$, where $\gamma^{\prime}$ is any other loop in $X^{0}$ which is homotopic to $\gamma$. Since $\pi_{1}(X^{0})\to\pi_{1}(X)$ is surjective with kernel generated by $\delta_{i}$’s, $\exp(2\pi\mathrm{i}\int_{\delta_{i}}\phi_{j})=1$ for all $i$, and $\phi_{j}$ is $d$-closed, the expression $\exp(2\pi\mathrm{i}\int_{\gamma}\phi_{j})$ is well-defined. Hence, we have (ii) by the representation (1).
(3) By $t_{j}(\gamma x)=\exp(2\pi\mathrm{i}\int_{\gamma}\phi_{j})t_{j}(x)$ for all $\gamma\in\pi_{1}(X)$, we have $t_{j}$ is a holomorphic nowhere vanishing section of $p^{*}L_{j}$ over $X^{0}$. Let $\{x_{i}\}$ be a coordinate system on an open subset $U$ of $X$ such that $D=(x_{1}x_{2}\cdots x_{c}=0)$. Then
@@
Hence $t_{j}$ is a meromorphic, and so rational (by Serre’s GAGA) section of $p^{*}L_{j}$ and $\mathrm{div}(t_{j})=\sum_{i}a_{ij}D_{i}$. ∎
Let $(z_{1},\cdots,z_{d})$ be the inhomogeneous fiber coordinates of the $(\mathbb{P}^{1})^{d}$-bundle $\mathbb{P}_{A}\to A$, the rational map $X\dashrightarrow\mathbb{P}_{A}$ over $A$ is now given by
@@
Let $H:=\mathbb{P}_{A}\setminus G=\sum_{i}((z_{i}=0)+(z_{i}=\infty))$. The pair $(\mathbb{P}_{A},H)$ is log smooth and $K_{\mathbb{P}_{A}}+H\sim 0$.
Example 10.
@@
(1) Let $X=\mathbb{P}^{2}$. If $D_{X}=$ 3 lines, then $q(X,D_{X})=3$. If $D_{X}=$ an elliptic curve, then $q(X,D_{X})=0$. If $D_{X}=Q+L$, where $L=(l=0)$ is a line and $Q=(q=0)$ is a conic, then $q(X,D_{X})=1$, $H^{0}(\Omega^{1}_{X}(\log D_{X}))=\mathbb{C}\cdot d\log\frac{l^{2}}{q}$ and the quasi-Albanese morphism is $X\setminus D\to\mathbb{C}^{*},x\mapsto\frac{l^{2}}{q}$.
(2) Let $X=\mathbb{P}^{1}_{z_{1}}\times\mathbb{P}^{1}_{z_{2}}$ and $D_{X}=F_{1}+F_{2}+\Delta$ where $F_{1}=(z_{1}=0)$, $F_{2}=(z_{2}=0)$, $\Delta=(z_{1}=z_{2})$. Then $H^{0}(\Omega_{X}(\log D_{X}))=\mathbb{C}\cdot d\log(1/z_{1}-1/z_{2})$ and the quasi-Albanese morphism is $X\setminus D\to\mathbb{C}^{*},(z_{1},z_{2})\mapsto 1/z_{1}-1/z_{2}$.
@@
Remark 11.
Let $X$ be a (possibly) singular variety. Let $\pi:X^{\prime}\to X$ be a resolution and let $f^{\prime}:X^{\prime}\to G$ be the quasi-Albanese morphism of $X^{\prime}$, then the quasi-Albanese map of $X$ is defined as the rational map $f^{\prime}\circ\pi^{-1}$. Assume that $X$ has klt singularities. Then the fibers of $\pi$ are rationally chain connected by Hacon-McKernan [HM07]. Since $G$ contains no rational curves, $f^{\prime}$ factors through $\pi$ by the rigidity Lemma (cf.[Deb01, Lemma 1.15]). The statement in [Deb01, Lemma 1.15] requires $f^{\prime}$ to be proper, but the proof only used that the image of $X^{\prime}\to X\times G$ is proper over $X$, which is guaranteed by [Har77, Chapter II Exercise 4.4].
Proposition 12.
Let $(X,D_{X})$ be a $\mathbb{Q}$-factorial dlt pair with $D_{X}$ reduced and $\kappa(K_{X}+D_{X})=0$. Then the quasi-Albanese morphism $X\setminus D_{X}\to G$ is dominant and has irreducible general fibers.
@@
Let $\pi:Y\to X$ be a log resolution of $(X,D_{X})$ and let $D_{Y}:=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$. Then $q(Y,D_{Y})=q(Y,\pi^{-1}D_{X})=q(X,D_{X})$ and $\kappa(K_{Y}+D_{Y})=0$ by Proposition 5. By the universal property of quasi-Albanese morphisms, we have the following commutative diagram.
@@
where $g$, $g^{\prime}$ and $f$ are quasi-Albanese morphisms. Since the injective homomorphism $H^{0}(\Omega^{1}_{Y}(\log\pi^{-1}D_{X}))\to H^{0}(\Omega_{Y}^{1}(\log D_{Y}))$ is an isomorphism, we have $H^{1}(Y\setminus D_{Y},\mathbb{C})=H^{1}(O_{Y})\oplus H^{0}(\Omega_{Y}(\log\pi^{-1}D_{X}))\to H^{1}(Y\setminus\pi^{-1}D_{X},\mathbb{C})=H^{1}(O_{Y})\oplus H^{0}(\Omega_{Y}^{1}(\log D_{Y}))$ is an isomorphism by Deligne’s theory of mixed Hodge structures. Hence the surjective morphism $H_{1}(Y\setminus D_{Y},\mathbb{Z})/{\mathrm{torsion}}\to H_{1}(Y\setminus\pi^{-1}D_{X},\mathbb{Z})/{\mathrm{torsion}}$ is also an isomorphism, so $G^{\prime}\to G$ is an isomorphism.
Since $\kappa(K_{Y}+D_{Y})=0$, we have $g^{\prime}$ is dominant and has irreducible general fibers by Theorem 1. Hence, $g$, $f$ are dominant and have irreducible general fibers. ∎
@@ 2.3. Subadditivity and positivity
Theorem 13.
Let $f:X\to Y$ be a fibration between smooth projective varieties. Let $D_{X}$ and $D_{Y}$ be reduced snc divisors on $X$ and $Y$ respectively. Denote by $F$ a general fiber of $f$. Assume that $f^{-1}D_{Y}\subset D_{X}$. Then the following holds.
(1)[Mae86, Corollary 2] Let $B$ be a big $\mathbb{Q}$-divisor on $Y$, then
$\kappa(K_{X}+D_{X}-f^{*}(K_{Y}+D_{Y})+f^{*}B)\geq\kappa(K_{F}+{D_{X}}|_{F})+\dim Y.$
(2) If $\kappa(K_{F}+{D_{X}}|_{F})\geq 0$, then $K_{X}+D_{X}-f^{*}(K_{Y}+D_{Y})$ is pseudoeffective.
@@
Since the $\mathbb{Q}$-divisor $B$ in (1) can be sufficiently small, (2) follows.
Let $X^{0}$ be a smooth quasi-projective variety and let $X^{0}\subset X$ be a compactification such that $D_{X}:=X\setminus X^{0}$ is a simple normal crossing divisor. By the same reasoning as in the definition of $\overline{\kappa}(X^{0})$ (cf. [Fuj24, Lemma 2.3]), we can define the numerical logarithmic Kodaira dimension as $\overline{\nu}(X^{0}):=\nu(K_{X}+D_{X})$. By Fujino’s subadditivity 13(3), we have the following.
Lemma 14.
Let $f:X\to Y$ be a dominant morphism with irreducible general fibers between smooth quasi-projective varieties. Let $F$ be a general fiber of $f$. Then $\overline{\nu}(X)\geq\overline{\nu}(F)+\overline{\kappa}(Y)$.

### v2

@@ 2.2. Quasi-Albanese morphisms for smooth varieties
We recall the constructions of quasi-Albanese morphisms in [iitaka1976logarithmic, fujino2024quasi]. Let $(X,D_{X}=\sum_{i=1}^{m}D_{i})$ be a projective log smooth pair with $D_{X}$ reduced. Where $D_{i}$’s are prime divisors. Let $X^{0}=X\setminus D_{X}$ and let $\iota:X^{0}\to X$ be the open embedding. Let $p:X\to A$ be the Albanese morphism. Set $q=q(X)$, $\overline{q}=q(X,D_{X})$, and $d=\overline{q}-q$.
Let $\eta_{i}$, $1\leq i\leq d$, be a basis of the free part of $\ker(\iota_{*}:H_{1}(X^{0},\mathbb{Z})\to H_{1}(X,\mathbb{Z}))$, and let $\xi_{i}\in H_{1}(X^{0},\mathbb{Z})$, $1\leq i\leq 2q$, be such that the $\eta_{i}$ and $\xi_{i}$ form a basis of the free part of $H_{1}(X^{0},\mathbb{Z})$.
Let $\omega_{i}$, $1\leq i\leq q$, be a basis of $H^{0}(\Omega_{X}^{1})$, and let $\phi_{i}\in H^{0}(\Omega_{X}^{1}(\log D))$, $1\leq i\leq d$, be such that the $\omega_{i}$ and $\phi_{i}$ form a basis of $H^{0}(\Omega_{X}^{1}(\log D_{X}))$, and satisfy $\int_{\eta_{i}}\omega_{j}=0$ and $\int_{\eta_{i}}\phi_{j}=\delta_{ij}$ (the Kronecker delta).
@@
$A_{i}=(\int_{\xi_{i}}\omega_{1},\cdots,\int_{\xi_{i}}\omega_{q},\int_{\xi_{i}}\phi_{1},\cdots,\int_{\xi_{i}}\phi_{d})=:(\hat{A}_{i},C_{i}),1\leq i\leq 2q,\text{ and }$
$B_{i}=(\int_{\eta_{i}}\omega_{1},\cdots,\int_{\eta_{i}}\omega_{q},\int_{\eta_{i}}\phi_{1},\cdots,\int_{\eta_{i}}\phi_{d})=:(0,\hat{B}_{i}),1\leq i\leq d.$
@@
$G:=\frac{H^{0}(X,\Omega_{X}^{1}(\log D))^{\vee}}{H_{1}(X^{0},\mathbb{Z})}=\frac{\mathbb{C}^{\overline{q}}}{\sum_{i}\mathbb{Z}A_{i}+\sum_{j}\mathbb{Z}B_{j}}.$
@@
$0\to{\mathbb{C}^{d}}/{\sum_{j}\mathbb{Z}\hat{B}_{j}}\cong(\mathbb{C}^{*})^{d}\to G\to{\mathbb{C}^{q}}/{\sum_{i}\mathbb{Z}\hat{A}_{i}}\to 0,$
where ${\mathbb{C}^{q}}/{\sum_{i}\mathbb{Z}\hat{A}_{i}}\cong A$, the Albanese variety of $X$, and the isomorphism $\mathbb{C}^{d}/\sum_{j}\mathbb{Z}\hat{B}_{j}\cong(\mathbb{C}^{*})^{d}$ is given by
@@
Since $G$ is quasi-abelian, there are $d=\dim G-\dim A$ line bundles $L_{j}\in\text{Pic}^{0}(A)$ such that $G$ is the fiber product over $A$ of the total spaces of the $L_{j}$ with their zero sections removed (cf. [Vojta1996Integral, Section 2]). The following content in this subsection aims to explore these line bundles $L_{j}$. The explicit form of $L_{j}$, together with Lemma 3.7, helps us calculate examples.
Let $G^{\prime}\to\mathbb{C}^{q}$ be the pullback of the $(\mathbb{C}^{*})^{d}$-bundle $G\to A$ by the cover $\mathbb{C}^{q}\to A$. Then we have $G^{\prime}=\mathbb{C}^{\overline{q}}/\sum_{i}\mathbb{Z}B_{i}\cong\mathbb{C}^{q}\times(\mathbb{C}^{*})^{d}$ and
@@
Hence $L_{j}$ is the line bundle corresponding to the representation
(2.2.1) $\pi_{1}(A)\to\mathbb{C}^{*},\hat{A}_{i}\mapsto\exp(2\pi\mathrm{i}\int_{\xi_{i}}\phi_{j}),1\leq i\leq q.$
@@
(2) By the expression $\exp(2\pi\mathrm{i}\int_{\gamma}\phi_{j})$, we mean $\exp(2\pi\mathrm{i}\int_{\gamma^{\prime}}\phi_{j})$, where $\gamma^{\prime}$ is any other loop in $X^{0}$ which is homotopic to $\gamma$. Since $\pi_{1}(X^{0})\to\pi_{1}(X)$ is surjective with kernel generated by $\delta_{i}$’s, $\exp(2\pi\mathrm{i}\int_{\delta_{i}}\phi_{j})=1$ for all $i$, and $\phi_{j}$ is $d$-closed, the expression $\exp(2\pi\mathrm{i}\int_{\gamma}\phi_{j})$ is well-defined. Hence, we have (2) by the representation (2.2.1).
(3) Since $t_{j}(\gamma x)=\exp(2\pi\mathrm{i}\int_{\gamma}\phi_{j})t_{j}(x)$ for all $\gamma\in\pi_{1}(X)$, we have $t_{j}$ is a holomorphic nowhere vanishing section of $p^{*}L_{j}$ over $X^{0}$. Let $\{x_{i}\}$ be a coordinate system on an analytic open subset $U$ of $X$ such that $D=(x_{1}x_{2}\cdots x_{c}=0)$. Then
@@
Hence $t_{j}$ is a meromorphic, and so a rational (by Serre’s GAGA) section of $p^{*}L_{j}$ and $\mathrm{div}(t_{j})=\sum_{i}a_{ij}D_{i}$. ∎
Let $(\mathbb{P}_{A},H)$ be the canonical compactification of $G$. Let $(z_{1},\cdots,z_{d})$ be the inhomogeneous fiber coordinates of the $(\mathbb{P}^{1})^{d}$-bundle $\mathbb{P}_{A}\to A$, the rational map $X\dashrightarrow\mathbb{P}_{A}$ over $A$ is now given by
@@
Example 2.5.
@@
(1) Let $X=\mathbb{P}^{2}$. If $D_{X}=$ 3 lines, then $q(X,D_{X})=3$, and $X\setminus D_{X}=(\mathbb{C}^{*})^{2}$ itself is a quasi-abelian variety. If $D_{X}=$ an elliptic curve, then $q(X,D_{X})=0$. If $D_{X}=Q+L$, where $L=(l=0)$ is a line and $Q=(q=0)$ is a conic, then $q(X,D_{X})=1$, $H^{0}(\Omega^{1}_{X}(\log D_{X}))=\mathbb{C}\cdot d\log\frac{l^{2}}{q}$ and the quasi-Albanese morphism is $X\setminus D\to\mathbb{C}^{*},x\mapsto\frac{l^{2}}{q}$. The rational map $X\dashrightarrow\mathbb{P}^{1}$, $x\mapsto[l^{2}:q]$ is not defined at $Q\cap L$.
(2) Let $X=\mathbb{P}^{1}_{z_{1}}\times\mathbb{P}^{1}_{z_{2}}$ and $D_{X}=F_{1}+F_{2}+\Delta$ where $F_{1}=(z_{1}=0)$, $F_{2}=(z_{2}=0)$, $\Delta=(z_{1}=z_{2})$. Then $H^{0}(\Omega_{X}^{1}(\log D_{X}))=\mathbb{C}\cdot d\log(1/z_{1}-1/z_{2})$ and the quasi-Albanese morphism is $X\setminus D\to\mathbb{C}^{*},(z_{1},z_{2})\mapsto 1/z_{1}-1/z_{2}$. The rational map $X\dashrightarrow\mathbb{P}^{1}$, $(z_{1},z_{2})\mapsto 1/z_{1}-1/z_{2}$ is not defined at $(F_{1}+F_{2})\cap\Delta=(z_{1}=z_{2}=0)$.
@@
Lemma 2.6.
Let $X$ be a smooth projective variety. Let $D_{X}\leq D_{X}^{\prime}$ be reduced snc divisors. By the universal property for quasi-Albanese morphisms, we have the following commutative diagram.
$X\setminus D_{X}^{\prime}$$G^{\prime}$$X\setminus D_{X}$}G{G}
where the horizontal arrows are quasi-Albanese morphisms and $G^{\prime}\to G$ is the induced morphism. Assume that $q(X,D_{X})=q(X,D_{X}^{\prime})$. Then $G^{\prime}\to G$ is an isomorphism.
@@
Since the injective homomorphism
$H^{0}(X,\Omega_{X}^{1}(\log D_{X}))\to H^{0}(X,\Omega_{X}^{1}(\log D_{X}^{\prime}))$
is an isomorphism, we have $H^{1}(X\setminus D_{X},\mathbb{C})\to H^{1}(X\setminus D_{X}^{\prime},\mathbb{C})$ is an isomorphism by Deligne’s mixed Hodge structures:
$H^{1}(X\setminus D_{X},\mathbb{C})$$H^{1}(X\setminus D_{X}^{\prime},\mathbb{C})$$H^{1}(X,O_{X})\oplus H^{0}(X,\Omega_{X}^{1}(\log D_{X}))$$H^{1}(X,O_{X})\oplus H^{0}(X,\Omega_{X}^{1}(\log D_{X}^{\prime}))$$\scriptstyle{\cong}\scriptstyle{\cong}\scriptstyle{\cong}$
Hence
$G^{\prime}=\frac{H^{0}(X,\Omega_{X}^{1}(\log D_{X}^{\prime}))^{\vee}}{H_{1}(Z\setminus D_{X}^{\prime},\mathbb{Z})}\to G=\frac{H^{0}(X,\Omega_{X}^{1}(\log D_{X}))^{\vee}}{H_{1}(Z\setminus D_{X},\mathbb{Z})}$
is an isomorphism. ∎
2.3. Quasi-Albanese morphisms for klt varieties
Let $X$ be a (possibly) singular variety. Let $\pi:X^{\prime}\to X$ be a resolution and let $f^{\prime}:X^{\prime}\to G$ be the quasi-Albanese morphism of $X^{\prime}$, then the quasi-Albanese map of $X$ is defined as the rational map $f^{\prime}\circ\pi^{-1}$. Assume that $X$ has klt singularities. Then the fibers of $\pi$ are rationally chain connected by Hacon-McKernan [hacon2007shokurov]. Since $G$ contains no rational curves, $f^{\prime}$ factors through $\pi$ by the rigidity Lemma (cf. [KollarRationalCurves, 5.3 Proposition]). In other words, there is a morphism $f:X\to G$ such that $f\circ\pi=f^{\prime}$.
@@ 2.3. Quasi-Albanese morphisms for klt varieties
One checks that the morphism $f:X\to G$ given in this way satisfies the universal property as in Theorem 2.4.
Let $(X,D_{X})$ be a projective lc pair and let $\operatorname{nklt}=\operatorname{nklt}(X,D_{X})$ be its non-klt locus. Let $\alpha:X\setminus\operatorname{nklt}\to G$ be the quasi-Albanese morphism. Let $A$ be the compact part of $G$ and let $A^{\prime}:=(\text{Pic}^{0}(X))^{\vee}_{\text{red}}$ be the Albanese variety of $X$. Let $\pi:Z\to X$ be a log resolution such that $\pi^{-1}\operatorname{nklt}$ is a divisor. By our definition, $G$ is the quasi-Albanese variety of $Z\setminus\pi^{-1}\operatorname{nklt}$. Hence $A$ is the Albanese variety of $Z$ by Iitaka’s construction. By the universal property for Albanese morphisms, we have the following commutative diagram.
[…]
where $A\to A^{\prime}$ is the morphism induced by $Z\to X$. Unlike the log smooth case, in general, we have $A\not\cong A^{\prime}$. But if $(X,D_{X})$ is dlt, then by [hacon2007shokurov] and the rigidity Lemma, we do have $A\cong A^{\prime}$.
Lemma 2.7.
Let $(X,D_{X})$ be a projective lc pair. Let $G$ be the quasi-Albanese variety of $X\setminus\operatorname{nklt}(X,D_{X})$. Let $\pi:Z\to X$ be a log resolution such that $D_{Z}^{\prime}:=\pi^{-1}\operatorname{nklt}(X,D_{X})$ is a divisor. Let $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$. Note that $D_{Z}^{\prime}\leq\lfloor D_{Z}\rfloor$. By the universal property for quasi-Albanese morphisms, we have the following commutative diagram.
$Z\setminus\lfloor D_{Z}\rfloor$$G^{\prime}$$Z\setminus D_{Z}^{\prime}$}G{G}
where the horizontal arrows are quasi-Albanese morphisms and $G^{\prime}\to G$ is the induced morphism. Then $G^{\prime}\to G$ is an isomorphism. In particular, we have $\dim G=q(X,D_{X})$.
Proof.
By Lemma 2.2, we have $q(X,D_{X})=q(Z,D_{Z}^{\prime})=q(Z,D_{Z})$. Hence the Lemma follows from Lemma 2.6. ∎
@@ 2.4. Subadditivity and positivity
The following is a summary of slightly generalizations of the relevant theorems due to Kawamata [kawamata1981characterization], Maehara [maehara1986weak], Campana [Campana2004Orbifolds] and Fujino [Fujino2017notes], and it is one of the most important techniques used in this note.
Theorem 2.8.
Let $f:X\to Y$ be a fibration between smooth projective varieties. Let $D_{X}$ and $D_{Y}$ be divisors on $X$ and $Y$ respectively, such that $(X,D_{X})$ and $(Y,D_{Y})$ are log smooth lc. Assume that $D_{Y}$ is reduced. Denote by $F$ a very general fiber of $f$. Assume that
$f^{-1}D_{Y}\subset\lfloor D_{X}\rfloor.$
Then the following holds.
(1)(cf. [kawamata1981characterization, Theorem 30], [maehara1986weak, Corollary 1, Corollary 2], [Campana2004Orbifolds, Theorem 4.2] or [Fujino2017notes, Theorem 1.9]) Let $B$ be a big $\mathbb{Q}$-divisor on $Y$, then
$\kappa(K_{X}+D_{X}-f^{*}(K_{Y}+D_{Y})+f^{*}B)=\kappa(K_{F}+{D_{X}}|_{F})+\dim Y.$
In particular, if $K_{Y}+D_{Y}$ is big, then
$\kappa(K_{X}+D_{X})=\kappa(K_{F}+D_{X}|_{F})+\dim Y.$
(2) If $\kappa(K_{F}+{D_{X}}|_{F})\geq 0$, then $K_{X}+D_{X}-f^{*}(K_{Y}+D_{Y})$ is pseudo-effective.
@@
Proof.
Note that we do not require $D_{X}$ to be reduced here. Using [Fujino2017notes, Theorem 1.1], the proof of (1) is identical to that of [Fujino2017notes, Theorem 1.9]. Since the $\mathbb{Q}$-divisor $B$ in (1) can be sufficiently small, (2) follows. ∎
Let $(X^{0},D^{0})$ be a quasi-projective log smooth lc pair, and let $X^{0}\subset X$ be a compactification such that $(X,D+B)$ is log smooth, where $B=X\setminus X^{0}$ and $D$ is the closure of $D^{0}$ in $X$. We define the logarithmic Kodaira dimension $\overline{\kappa}$ (resp. numerical logarithmic Kodaira dimension $\overline{\nu}$) of $(X^{0},D^{0})$ as $\overline{\kappa}(X^{0},D^{0}):=\kappa(K_{X}+D+B)$ (resp. $\overline{\nu}(X^{0},D^{0}):=\nu(K_{X}+D+B)$). To show that $\overline{\kappa}$ and $\overline{\nu}$ are well defined, it suffices to show that any two such compactifications give the same $\overline{\kappa}$ and $\overline{\nu}$. Let $(X,D+B)$ and $(X^{\prime},D^{\prime}+B^{\prime})$ be two such compactifications. Let $X\xleftarrow{p}X^{\prime\prime}\xrightarrow{q}X^{\prime}$ be a common resolution, which is an isomorphism over $X^{0}$, and such that $D^{\prime\prime}+B^{\prime\prime}$ is snc, where $B^{\prime\prime}=X^{\prime\prime}\setminus X^{0}$ and $D^{\prime\prime}$ is the closure of $D^{0}$ in $X^{\prime\prime}$. Then $D^{\prime\prime}=p^{-1}_{*}D$ and $B^{\prime\prime}=p^{-1}_{*}B+\operatorname{Ex}(p)$. Hence
$K_{X^{\prime\prime}}+D^{\prime\prime}+B^{\prime\prime}=p^{*}(K_{X}+D+B)+\sum_{E\subset\operatorname{Ex}(p)}(a(E,X,D+B)+1)E.$
Since $(X,D+B)$ is lc, we have $\sum_{E\subset\operatorname{Ex}(p)}(a(E,X,D+B)+1)E\geq 0$. Hence $\kappa(K_{X}+D+B)=\kappa(K_{X^{\prime\prime}}+D^{\prime\prime}+B^{\prime\prime})=\kappa(K_{X^{\prime}}+D^{\prime}+B^{\prime})$ and $\nu(K_{X}+D+B)=\nu(K_{X^{\prime\prime}}+D^{\prime\prime}+B^{\prime\prime})=\nu(K_{X^{\prime}}+D^{\prime}+B^{\prime})$. Using these definitions, we can restate Theorem 2.8 (1) (3) as follows.
Lemma 2.9.
Let $f^{0}:X^{0}\to Y^{0}$ be a fibration between smooth quasi-projective varieties. Let $D^{0}$ be a divisor on $X^{0}$ such that $(X^{0},D^{0})$ is log smooth lc. Denote by $F^{0}$ a very general fiber of $f^{0}$. Then the following holds.
  1. (1)
    If $\overline{\kappa}(Y^{0})=\dim Y^{0}$, then $\overline{\kappa}(X^{0},D^{0})=\overline{\kappa}(F^{0},D^{0}|_{F^{0}})+\dim Y^{0}$.
  2. (2)
    $\overline{\nu}(X^{0},D^{0})\geq\overline{\nu}(F^{0},D^{0}|_{F^{0}})+\overline{\kappa}(Y^{0})$.
Proof.
Let $X^{0}\subset X$ and $Y^{0}\subset Y$ be compactifications, such that $(X,D+B_{X})$, $(Y,B_{Y})$ are log smooth and there is a morphism $f:X\to Y$ extending $f^{0}$, where $D$ is the closure of $D^{0}$ in $X$, $B_{X}=X\setminus X^{0}$ and $B_{Y}=Y\setminus Y^{0}$. By construction, we have $f^{-1}B_{Y}\subset B_{X}\subset\lfloor D+B_{X}\rfloor$. Then (1), (2) follow from Theorem 2.8 (1), (3) respectively. ∎